OpenStudy (lena772):

(a) find an equation of the tangent line to the graph of f at the given point, (b) use a graphing utility to graph the function and its tangent line at the point, and (c) use the derivative feature of a graphing utility to confirm your results.

OpenStudy (lena772):

\[f(x)=x^2+2x+1 \]

OpenStudy (lena772):

(-3, 4)

ganeshie8 (ganeshie8):

you wanto find the tangent line thru (-3, 4)

ganeshie8 (ganeshie8):

you need to knw slope to write the equation of tangent, slope of tangent at x = -3, is \(f'(-3)\)

ganeshie8 (ganeshie8):

start by finding \(f'(3)\)

ganeshie8 (ganeshie8):

\(\large f'(3) = \lim \limits_{x \to 3} \frac{f(x) - f(3)}{x-3}\)

OpenStudy (lena772):

\[f(3)=3^2+2*3+\]

ganeshie8 (ganeshie8):

yes do it all the way in ur notes, and tell me wat u get for \(f'(3)\)

OpenStudy (lena772):

\[f(3)=3^2+2*3+1\]

OpenStudy (lena772):

\[f(3)=9+6+1, f(3)=16\]

ganeshie8 (ganeshie8):

yes, plug it above

OpenStudy (lena772):

\[f'(3)=\lim_{x \rightarrow 3}\frac{ f(x)-16 }{ x-3 } \]

ganeshie8 (ganeshie8):

yes, keep going

OpenStudy (lena772):

i'm stuck :(

OpenStudy (lena772):

wait no

ganeshie8 (ganeshie8):

:)

OpenStudy (lena772):

\[f'(3) = \lim_{x \rightarrow 3}\frac{ (x^3+2x+1)-(16) }{ x-3}\]

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

wat next

OpenStudy (lena772):

\[f'(x)=\lim_{x \rightarrow 3}\frac{ (x^3+2x-15) }{(x-3) }\]

ganeshie8 (ganeshie8):

factor numerator and see if it cancels the bottom illegal thingy :)

OpenStudy (lena772):

3-3=0.

ganeshie8 (ganeshie8):

hey, how come x^2 became x^3 :o

OpenStudy (lena772):

woops sorry, it should be x^2

ganeshie8 (ganeshie8):

\(\large f'(x)=\lim \limits_{x \rightarrow 3}\frac{ (x^{\color{red}{2}}+2x-15) }{(x-3) } \)

ganeshie8 (ganeshie8):

Yes :) try to factor numerator

OpenStudy (lena772):

(x(x+2))-15

ganeshie8 (ganeshie8):

no not like that

ganeshie8 (ganeshie8):

\(x^2+2x-15\) think of two numbers such that sum = 2 product = -15

ganeshie8 (ganeshie8):

5-3 = 2 5*-3 = -15 so, \(x^2+2x-15 = (x+5)(x-3)\)

OpenStudy (lena772):

oh yeah

ganeshie8 (ganeshie8):

plug them above, and tell me wat u get for final f'(3)

OpenStudy (lena772):

f'(x)=limx-->3 x+5

ganeshie8 (ganeshie8):

Yes, now that the bottom illegal thing is gone, you can take the limit now plug x = 3

ganeshie8 (ganeshie8):

f'(x)=limx-->3 x+5 = 3+5 = 8

OpenStudy (lena772):

f'(3)=8

ganeshie8 (ganeshie8):

Yes, f'(3) is the slope of given function at x = 3 so you knw the point and u knw the slope

ganeshie8 (ganeshie8):

you can write the equation of line in point-slope form

OpenStudy (lena772):

Should I be putting f'(x) in front of the lim or f'(3)

ganeshie8 (ganeshie8):

Very good question :)

ganeshie8 (ganeshie8):

since we're finding the derivative AT x = 3 we should put f'(3) oly

ganeshie8 (ganeshie8):

if you're finding derivative in general for ANY X, then we put f'(x) again, its just a notational thing

ganeshie8 (ganeshie8):

so, wats the equation of tangent line ?

OpenStudy (lena772):

y-4=8(x--3) y-4=8(x+3)

ganeshie8 (ganeshie8):

Awesome :)

OpenStudy (lena772):

So that is the answer for A?

ganeshie8 (ganeshie8):

yes

OpenStudy (lena772):

Ok for b can i just put that equation in GeoGebra to get the right graph

ganeshie8 (ganeshie8):

Yes ! put the equation for f(x) in geogebra

ganeshie8 (ganeshie8):

plot the point (-3, 4) also

OpenStudy (lena772):

What should I label the point?

ganeshie8 (ganeshie8):

A = (-3, 4)

ganeshie8 (ganeshie8):

or P = (-3, 4)

OpenStudy (lena772):

\[f(x)=x^2+2x+1 \]

ganeshie8 (ganeshie8):

anything wil do

OpenStudy (lena772):

Put that in geogebra?

ganeshie8 (ganeshie8):

yes ! plugin both funciton and point

ganeshie8 (ganeshie8):

and graph the tangent line which u got also : y-4=8(x+3)

ganeshie8 (ganeshie8):

graph below 3 :- 1) function f(x) = x^2+2x+1 2) point P = (-3, 4) 3) tangent y-4=8(x+3)

OpenStudy (lena772):

how do i use the derivative feature to check

ganeshie8 (ganeshie8):

looks good, let me open geogebra and try if we can plot the derivative function

OpenStudy (lena772):

will my teacher know if i checked it or not?

ganeshie8 (ganeshie8):

she wont, but if u have time... it is goot to confirm girly :)

OpenStudy (lena772):

lol don't worry about it i trust you

ganeshie8 (ganeshie8):

:) once i get to knw how to do it in geogebra, il let u knw...

OpenStudy (lena772):

thanks!